OpenStudy (anonymous):

write in standard form. Determine the value of a,b,c 7x^2-3x=0 a= b= c=

OpenStudy (anonymous):

Compare to ax^2+bx+c=0

OpenStudy (anonymous):

do the problem so i can understand it

OpenStudy (anonymous):

ax^2+bx+c=0 (7)x^2+(-3)x+(0)=0 a=7, b=-3 , c=0